OpenStudy (anonymous):

Prove tan^2x+sec^2x=1

OpenStudy (amistre64):

t^2 + 1 = s^2

OpenStudy (amistre64):

s^2 + c^2 = 1 ------------- c^2 t^2 + 1 = s^2

OpenStudy (anonymous):

"proof" let \[x=\frac{\pi}{4}\]

OpenStudy (anonymous):

\[ \tan^2x+\sec^2x=1 \] This is not true in our universe.

OpenStudy (anonymous):

then \[\tan^2(\frac{\pi}{4})+\sec^2(\frac{\pi}{4})\] \[=1^2+(\frac{2}{\sqrt{2}})^2\] \[=1+2=3\]

OpenStudy (amistre64):

but since we are part of a multiverse .... :)

OpenStudy (amistre64):

pbs rotted my brain!!

OpenStudy (anonymous):

pbr rotted mine

OpenStudy (amistre64):

pbjs was my next option for consideration lol